OpenStudy (anonymous):

a basketball player jumps up and is in the air for 0.7s. how high did she jump?

OpenStudy (timmyk14):

D = S * T You need to know the speed and the time for this. All you've given us is the time. Is there a speed?

OpenStudy (anonymous):

nahh. the question only mentions that much :\

OpenStudy (timmyk14):

Interesting...Well I'm sorry I can't answer this; unless this is a multiple choice question. In that case we could substitute and find the value of S, to see which distance would fit the equation.

OpenStudy (anonymous):

ouu. yahh. its a multiple choice ques. lol

OpenStudy (timmyk14):

Post them and I'll see if I can figure it out :)

OpenStudy (anonymous):

kk hold on

OpenStudy (anonymous):

a) 0.3m b) 0.6m c) 0.9m d) 1.2 m

OpenStudy (timmyk14):

Sorry the page messed up so I need to restart.

OpenStudy (timmyk14):

S = D/T S = 0.9/.7 S= 1.29 (rounded to the nearest hundredth) Plug this in to solve for D: D = S * T D = 1.29 * 0.7 D = 0.903 (Round this to the nearest hundredth and you have what?? 0.9!!) Get it?

OpenStudy (anonymous):

thankkss